# Jailbreak Antidote: Runtime Safety-Utility Balance via Sparse Representation Adjustment in Large Language Models

### v2

@@ 1 Introduction
Large language models (LLMs) have revolutionized natural language processing, demonstrating advanced cognitive abilities and significantly impacting various aspects of daily life. They excel in instruction understanding (Ouyang et al., 2022; Chung et al., 2024), summarization (Chung et al., 2024), and complex reasoning tasks (Kojima et al., 2022; Wang & Zhou, 2024). Applications built upon LLMs are widespread, enhancing efficiency and convenience in domains such as coding assistance (Roziere et al., 2023), medical diagnostics (Singhal et al., 2023), financial analysis (Li et al., 2023), and psychological counseling (Strachan et al., 2024; Xu et al., 2024). Given their pervasive use and profound social impact, ensuring the safety and utility of LLMs has become critically important.
@@ 2 Related Work
As LLMs become increasingly prevalent, they have become targets for jailbreak attacks—adversarial prompts designed to bypass safety mechanisms and induce models to generate harmful or disallowed content (Jin et al., 2024). Early attacks exploited simple manipulations like role-playing scenarios or specific prompts to trick models into violating safety guidelines (Wei et al., 2024). As safety alignment techniques improved, attackers developed more sophisticated methods, including gradient-based approaches that generate adversarial suffixes (Zou et al., 2023b), genetic algorithms to produce stealthy prompts (Liu et al., 2024), and black-box attacks that iteratively refine prompts without access to internal parameters (Chao et al., 2023). Other techniques involve crafting adversarial paraphrases (Zeng et al., 2024) or exploiting unconventional inputs like ciphered text (Yuan et al., 2024) and past tense formulations (Andriushchenko & Flammarion, 2024). These diverse and evolving attacks highlight the urgent need for robust defenses to maintain LLM safety and reliability.
@@
Existing defense strategies include prompt engineering, and safety fine-tuning. Detection-based approaches aim to identify and block adversarial prompts using techniques like perplexity filtering (Alon & Kamfonas, 2023; Jain et al., 2023), but sophisticated attacks with semantic-level prompts (Samvelyan et al., 2024; Paulus et al., 2024) often evade detection. Prompt engineering modifies prompts or model responses to reinforce safety, employing self-reminders (Xie et al., 2023) or leveraging in-context learning (Wei et al., 2023), but introduces computational overhead and inference latency (Agarwal et al., 2023), negatively affecting user experience. Safety alignment methods, such as Reinforcement Learning from Human Feedback (RLHF) (Bai et al., 2022) and Safe RLHF (Dai et al., 2024), retrain models on curated datasets but require significant resources and lack flexibility for real-time adjustments. Moreover, these methods may degrade model utility by being overly restrictive, leading to refusals of benign queries. We aim to address these limitations by proposing a defense mechanism that operates during inference without modifying input prompts or requiring retraining, enabling real-time safety adjustments while preserving model utility.
@@
Mechanistic interpretability seeks to reverse-engineer models by analyzing their internal representations (Elhage et al., 2021; Nanda et al., 2023). Prior research has explored how models process tasks like modular arithmetic and factual recall (Meng et al., 2022), focusing on interpretability rather than behavior control. Inspired by representation engineering (Zou et al., 2023a) and latent space steering (Liu et al., 2023), our work shifts focus to manipulating internal activations to adjust model behavior during inference.
A key finding of our approach is that safety representations in LLMs are sparsely distributed, allowing us to control the model’s safety preferences by adjusting only about $5\%$ of its internal state. This targeted manipulation contrasts with previous studies that often focus on higher-level structures like layers or attention heads (Halawi et al., 2024). While earlier works have primarily emphasized interpretability, we demonstrate that small-scale adjustments can directly influence LLM safety without the need for prompt modifications or retraining. This allows for efficient real-time adjustments, striking a balance between safety and utility with minimal performance impact.
@@ 3 Preliminaries
A successful jailbreak attack occurs when the model accepts a harmful prompt and generates a harmful response, i.e., when $\mathcal{J}(S_{0},R^{\prime})=1$, where $\mathcal{J}$ is a judge function. Various methods can implement the judge function, such as prefix matching (Zou et al., 2023b), LLM-based evaluations (Qi et al., 2024; Chao et al., 2023), or human annotations (Wei et al., 2023).
@@ 4.1 Identifying and Leveraging the Safety Direction
The first principal component $\mathbf{d}_{\text{safe}}^{l}$ captures the direction of maximum variance between benign and harmful prompts. In Figure 2 (a), the arrows represent the safety direction $\mathbf{d}_{\text{safe}}^{l}$ at different layers. We compute $\mathbf{d}_{\text{safe}}^{l}$ using only benign and harmful prompts, without including any adversarial attack prompts, to ensure generalization and avoid data leakage.
@@ 4.2 Sparsity in the Safety Representation
An important insight from our analysis is that the elements of $\mathbf{d}_{\text{safe}}^{l}$ exhibit a long-tail distribution, as shown in Figure 2 (b). This suggests that only a small subset of dimensions significantly contribute to the safety distinction, indicating that the safety representation in LLMs is sparse. To leverage this sparsity, we create a mask $\mathbf{m}^{l}\in\{0,1\}^{d}$ that retains only the top $k\%$ of dimensions with the largest absolute values in $\mathbf{d}_{\text{safe}}^{l}$:
@@ 5.2 Results and Analysis
We evaluated the overhead introduced by different defense methods by measuring the number of defense tokens required, which correlates with resource consumption and inference latency. Figure 4 shows scatter plots of Defense Tokens versus DSR for different defense methods across various models.
Figure 4: Defense Tokens versus DSR for different defense methods across various models. Each point represents a defense method, with the x-axis showing the number of defense tokens and the y-axis showing the DSR.
As shown in Figure 4, Jailbreak Antidote requires no additional prompt tokens, which means it introduces no overhead in terms of prompt length. This is because our method works by directly adjusting the internal states rather than modifying the input prompt. In contrast, methods like SemanticSmoothLLM and SmoothLLM rely on a significant number of defense tokens, leading to increased computational costs and inference delays. Despite this higher token consumption, some of these methods still achieve lower DSRs compared to our approach, indicating that their defense performance is not as effective relative to the computational overhead they introduce.
Our method consistently delivers high DSR with minimal inference overhead across all models, showcasing both its effectiveness and efficiency. These findings demonstrate that Jailbreak Antidote not only improves defense performance but also ensures inference efficiency, making it a practical solution for real-world applications.
@@ 5.4 Conclusion
In this work, we introduced Jailbreak Antidote, a method to enhance the safety of large language models (LLMs) by adjusting their internal states in real-time. By leveraging the sparsity of safety-related representations, we selectively modify a small subset of neurons (around 5%) to balance safety and utility without adding computational overhead. Extensive experiments across models from 2B to 72B parameters demonstrate that Jailbreak Antidote outperforms existing defenses in terms of Defense Success Rate (DSR) while maintaining high performance on benign tasks. This runtime flexibility offers a significant advantage for real-world applications, where safety needs may vary, but retraining or prompt modifications are impractical.
@@ A.1.1 Datasets
For evaluating safety and defense effectiveness, we used JailbreakBench (Chao et al., 2024), an open-source robustness benchmark for jailbreaking LLMs. JailbreakBench comprises $200$ distinct prompts, including $100$ benign and $100$ misuse prompts, curated with reference to OpenAI’s usage policies. We specifically used the $100$ misuse prompts as targets for jailbreak attacks to assess the robustness of different defense methods.
@@ A.1.2 Attack Methods
For all attacks, we used the successful prompts provided in the respective studies, ensuring consistency and reproducibility. We applied these attacks across different models to evaluate their robustness comprehensively.
@@ A.2.3 Impact of Token Position on Safety Representation
We investigated how the position of tokens affects the safety representation by computing the dot product between the hidden states of each token and the safety direction $\mathbf{d}_{\text{safe}}$. Figure A.3 shows heatmaps of these dot products across different layers and token positions for both benign and harmful prompts. The results highlight that the hidden state of the last token provides the most significant distinction between benign and harmful prompts, justifying our focus on adjusting the internal state at the last token position.
Figure A.3: Visualization of the dot product between hidden states and $\mathbf{d}_{\text{safe}}$ across layers and token positions. The last token (rightmost column) shows the most significant differentiation between benign and harmful prompts.

### v4

@@ 1 Introduction
Large language models (LLMs) have revolutionized natural language processing, demonstrating advanced cognitive abilities and significantly impacting various aspects of daily life. They excel in instruction understanding (Ouyang et al., 2022; Chung et al., 2024), summarization (Chung et al., 2024), and complex reasoning tasks (Kojima et al., 2022; Wang & Zhou, 2024). Applications built upon LLMs are widespread, enhancing efficiency and convenience in domains such as coding assistance (Roziere et al., 2023), medical diagnostics (Singhal et al., 2023), financial analysis (Li et al., 2023b), and psychological counseling (Strachan et al., 2024; Xu et al., 2024a). Given their pervasive use and profound social impact, ensuring the safety and utility of LLMs has become critically important Yi et al. (2024).
@@ 2 Related Work
As LLMs become increasingly prevalent, they have become targets for jailbreak attacks—adversarial prompts designed to bypass safety mechanisms and induce models to generate harmful or disallowed content (Jin et al., 2024). Early attacks exploited simple manipulations like role-playing scenarios or specific prompts to trick models into violating safety guidelines (Wei et al., 2024a). As safety alignment techniques improved, attackers developed more sophisticated methods, including gradient-based approaches that generate adversarial suffixes (Zou et al., 2023b), genetic algorithms to produce stealthy prompts (Liu et al., 2024), and black-box attacks that iteratively refine prompts without access to internal parameters (Chao et al., 2023). Other techniques involve crafting adversarial paraphrases (Zeng et al., 2024) or exploiting unconventional inputs like ciphered text (Yuan et al., 2024) and past tense formulations (Andriushchenko & Flammarion, 2024). These diverse and evolving attacks highlight the urgent need for robust defenses to maintain LLM safety and reliability.
@@
Existing defense strategies include prompt engineering, and safety fine-tuning. Detection-based approaches aim to identify and block adversarial prompts using techniques like perplexity filtering (Alon & Kamfonas, 2023; Jain et al., 2023), but sophisticated attacks with semantic-level prompts (Samvelyan et al., 2024; Paulus et al., 2024; Li et al., 2023a) often evade detection. Prompt engineering modifies prompts or model responses to reinforce safety, employing self-reminders (Xie et al., 2023) or leveraging in-context learning (Wei et al., 2023), but introduces computational overhead and inference latency (Agarwal et al., 2023), negatively affecting user experience. Safety alignment methods, such as Reinforcement Learning from Human Feedback (RLHF) (Bai et al., 2022) and Safe RLHF (Dai et al., 2024), retrain models on curated datasets but require significant resources and lack flexibility for real-time adjustments. Some approaches also defend against attacks by controlling the decoding process (Xu et al., 2024b), but require reference models and additional inference-time costs. Moreover, these methods may degrade model utility by being overly restrictive, leading to refusals of benign queries. We aim to address these limitations by proposing a defense mechanism that operates during inference without modifying input prompts or requiring retraining, enabling real-time safety adjustments while preserving model utility.
@@
Mechanistic interpretability seeks to reverse-engineer models by analyzing their internal representations (Elhage et al., 2021; Nanda et al., 2023). Prior research has explored how models process tasks like modular arithmetic and factual recall (Meng et al., 2022), focusing on interpretability rather than behavior control. Inspired by representation engineering (Zou et al., 2023a) and latent space steering  (Liu et al., 2023; Wei et al., 2024b; Turner et al., 2023), our work shifts focus to manipulating internal activations to adjust model behavior during inference.
Our key finding is that safety-related representations in LLMs are sparsely distributed, enabling effective control of the model’s safety preferences by modifying only about $5\%$ of its internal activations. This sparsity-based approach contrasts with previous studies that often target broader structures like layers or attention heads (Halawi et al., 2024). By demonstrating that small-scale, targeted adjustments can directly influence LLM safety, we move beyond interpretability to practical behavior control. Our method requires no prompt modifications or retraining, enabling efficient, real-time safety adjustments with minimal impact on utility and performance.
@@ 3 Preliminaries
A successful jailbreak attack occurs when the model accepts a harmful prompt and generates a harmful response, i.e., when $\mathcal{J}(S_{0},R^{\prime})=1$, where $\mathcal{J}$ is a judge function. Various methods can implement the judge function, such as prefix matching (Zou et al., 2023b), LLM-based evaluations (Qi et al., 2024b; Chao et al., 2023), or human annotations (Wei et al., 2023).
@@ 4.1 Identifying and Leveraging the Safety Direction
We then perform eigenvalue decomposition of the covariance matrix $\mathbf{C}^{l}$, which can be expressed as:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{C}^{l}=\mathbf{U}^{l}\mathbf{\Lambda}^{l}(\mathbf{U}^{l})^{\top},}$ (4)
where $\mathbf{U}^{l}=[\mathbf{u}_{1}^{l},\mathbf{u}_{2}^{l},\dots,\mathbf{u}_{d}^{l}]\in\mathbb{R}^{d\times d}$ is the orthogonal matrix whose columns $\mathbf{u}_{i}^{l}$ are the eigenvectors of $\mathbf{C}^{l}$, and $\mathbf{\Lambda}^{l}=\text{diag}(\lambda_{1}^{l},\lambda_{2}^{l},\dots,\lambda_{d}^{l})$ is the diagonal matrix of eigenvalues $\lambda_{1}^{l}\geq\lambda_{2}^{l}\geq\dots\geq\lambda_{d}^{l}$, representing the variance along the corresponding eigenvectors. The principal component $\mathbf{d}_{\text{safe}}^{l}$ is defined as the eigenvector $\mathbf{u}_{1}^{l}$ associated with the largest eigenvalue $\lambda_{1}^{l}$:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{d}_{\text{safe}}^{l}=\mathbf{u}_{1}^{l}.}$ (5)
The first principal component $\mathbf{d}_{\text{safe}}^{l}$ captures the direction of maximum variance between benign and harmful prompts. In Figure 2 (a), the arrows represent the safety direction $\mathbf{d}_{\text{safe}}^{l}$ at different layers. We compute $\mathbf{d}_{\text{safe}}^{l}$ using only benign and harmful prompts, without including any adversarial attack prompts, to ensure generalization and avoid data leakage. The points in the figure are visualized using t-SNE to illustrate the separation between benign and harmful prompts.
@@ 4.2 Sparsity in the Safety Representation
An important insight from our analysis is that the elements of $\mathbf{d}_{\text{safe}}^{l}$ exhibit a long-tail distribution, as shown in Figure 2 (b). This suggests that only a small subset of dimensions significantly contribute to the safety distinction, indicating that the safety representation in LLMs is sparse. Figure 2 (b) further emphasizes this sparsity by illustrating the dominance of a few components across layers. To leverage this sparsity, we create a mask $\mathbf{m}^{l}\in\{0,1\}^{d}$ that retains only the top $k\%$ of dimensions with the largest absolute values in $\mathbf{d}_{\text{safe}}^{l}$.
@@ 5.2 Results and Analysis
Comparison with Safety Alignment Defenses
To provide a more comprehensive evaluation, we include comparisons with safety alignment defenses, such as preference-based fine-tuning approaches (Qi et al., 2024a; Zou et al., 2024). For aligned models, AlpacaEval Win Rate is computed relative to their corresponding original models (e.g., Llama-3-8B-it-RR relative to Llama-3-8B-it). Our results show that Jailbreak Antidote not only achieves higher DSR compared to fine-tuned models, but also balances safety and utility more effectively. Furthermore, Jailbreak Antidote is fully compatible with fine-tuned models, enhancing their safety even further. This demonstrates the robustness and flexibility of our approach, which provides strong standalone performance while synergizing effectively with state-of-the-art alignment techniques.
Table 2: Comparison of DSR and Win Rate across different defense methods, including safety alignment defenses. The best, second and third scores are highlighted.
Model Safety -Utility
Baseline
In-Context Learning
Paraphrase
Perplexity Filter
Self Reminder
Semantic Smooth
Smooth LLM
Jailbreak Antidote
Llama-3-8B-it DSR $\uparrow$ 68.9 71.7 79.0 67.9 78.9 88.1 84.2 99.4
Win Rate $\uparrow$ 50.0 38.9 35.5 52.2 39.4 31.8 32.4 53.0
Llama-3-8B-it-RR DSR $\uparrow$ 77.0 77.2 91.1 77.3 80.5 92.2 95.6 99.6
(Zou et al., 2024) Win Rate $\uparrow$ 51.6 36.4 32.6 50.6 42.2 35.6 31.5 53.5
Gemma-2-9B-it DSR $\uparrow$ 54.5 56.7 75.8 55.1 63.1 79.4 46.5 78.1
Win Rate $\uparrow$ 50.0 38.6 31.2 51.0 42.5 33.9 32.4 47.4
Gemma-2-9B-it-DSA DSR $\uparrow$ 64.2 65.5 81.1 63.9 69.5 83.9 51.7 83.6
(Qi et al., 2024a) Win Rate $\uparrow$ 48.6 36.4 28.6 48.9 39.0 34.7 32.8 48.6
@@
We evaluated the overhead introduced by different defense methods by measuring the runtime per query, which represents the average time taken to process a single query during inference. This metric provides a practical and interpretable measure of efficiency for real-world applications, as it directly reflects the time required for generating responses. Figure 4 presents scatter plots of Runtime per Query versus DSR for various defense methods across different models.
Figure 4: Runtime per Query versus DSR for different defense methods across various models. Each point represents a defense method, with the x-axis showing the average runtime per query (seconds) and the y-axis showing the DSR.
As shown in Figure 4, Jailbreak Antidote achieves the shortest runtime per query across all models, highlighting its efficiency advantage. This is because our method works by directly adjusting the internal states rather than introducing additional tokens or modifying the input prompt, thus minimizing computational overhead. In contrast, methods like SemanticSmoothLLM and SmoothLLM result in significantly higher query runtimes due to their reliance on a substantial number of additional defense tokens, which increase computational cost and user-perceived delays. Despite their longer query runtimes, these methods achieve lower DSRs compared to our approach, indicating that their defense performance is less effective relative to the computational overhead they introduce.
To provide a hardware-agnostic perspective, we also include an analysis based on the number of defense tokens required in Appendix A.2.11. This complementary analysis correlates strongly with resource consumption and inference latency, particularly the Time to First Token (TTFT), and provides a consistent basis for comparison across different hardware platforms and inference engines.
@@ 5.4 Conclusion
In this work, we introduced Jailbreak Antidote, a method to enhance the safety of large language models (LLMs) by adjusting their internal states in real-time. Leveraging the sparsity of safety-related representations, we selectively modify a small subset of neurons to balance safety and utility without adding computational overhead. Extensive experiments across models from 2B to 72B parameters demonstrate that Jailbreak Antidote outperforms existing defenses in terms of Defense Success Rate (DSR) while maintaining high performance on benign tasks. However, our method also reveals a potential vulnerability: if an attacker manipulates the scaling factor $\alpha$ to negative values, they can shift the internal states toward unsafe directions, reducing the model’s safety. This dual nature underscores the challenges in defending against highly adaptive adversaries who might exploit such mechanisms.
@@
Acknowledgments
This work was supported in part by the Beijing Major Science and Technology Project under Contract No.Z241100001324005.
@@ A.1.1 Datasets
For evaluating safety and defense effectiveness, we used JailbreakBench (Mazeika et al., 2024), an open-source robustness benchmark for jailbreaking LLMs. JailbreakBench comprises $200$ distinct prompts, including $100$ benign and $100$ misuse prompts, curated with reference to OpenAI’s usage policies. We specifically used the $100$ misuse prompts as targets for jailbreak attacks to assess the robustness of different defense methods.
@@
To address concerns about similarity between the dataset used to generate safety directions  Phan (2023) and the evaluation dataset (JailbreakBench (Mazeika et al., 2024)), we conducted a similarity analysis using multiple metrics, summarized in Table A.1.
Table A.1: Similarity metrics between Phan (2023) and JailbreakBench (Mazeika et al., 2024).
Metric Value
TF-IDF Cosine Similarity 0.038
1-gram & 2-gram Jensen-Shannon Distance 0.547
BERT Cosine Similarity 0.768
The low TF-IDF similarity (0.038) and moderate Jensen-Shannon distance (0.547) indicate clear differences between the datasets. The BERT Cosine Similarity (0.768) is also lower than the similarity between benign and harmful subsets of JailbreakBench (0.840), confirming sufficient distinction between the datasets.
@@ A.1.2 Attack Methods
AutoDAN Attack
AutoDAN by Liu et al. (2024) uses a hierarchical genetic algorithm to generate stealthy, semantically meaningful jailbreak prompts, achieving strong transferability and bypassing perplexity-based defenses.
For all attacks, we used the successful prompts provided in the respective studies, such as those from JailbreakBench Chao et al. (2024) [^6] https://github.com/JailbreakBench/artifacts/tree/main/attack-artifacts, ensuring consistency and reproducibility. We applied these attacks across different models to evaluate their robustness comprehensively. This static testing approach allowed us to efficiently explore the large space of attack methods, defense mechanisms, and model combinations, balancing computational feasibility with experimental coverage.
To address the potential limitations of static attacks, we also incorporated adaptive attacks into our evaluation, as shown in Table A.4. Specifically, we utilized GCG (Zou et al., 2023b), PAIR (Chao et al., 2023), and AutoDAN (Liu et al., 2024) as representative adaptive attack methods. These methods dynamically adjust their strategies to target specific defense mechanisms, providing a stricter and more nuanced assessment of the robustness of our proposed method. For these experiments, we fixed the settings of Jailbreak Antidote and applied the adaptive attacks to evaluate its performance under more challenging scenarios.
@@ A.2.3 Impact of Token Position on Safety Representation
We investigated how the position of tokens affects the safety representation by computing the dot product between the hidden states of each token and the safety direction $\mathbf{d}_{\text{safe}}$. Figures A.3, A.4, A.5, and A.6 show heatmaps of these dot products across different layers and token positions for both benign and harmful prompts. The results highlight that the hidden state of the last token provides the most significant distinction between benign and harmful prompts, justifying our focus on adjusting the internal state at the last token position.
Figure A.3: Visualization of the dot product between hidden states and $\mathbf{d}_{\text{safe}}$ across layers and token positions on Llama-3-8B-it. The last token (rightmost column) shows the most significant differentiation between benign and harmful prompts.
Figure A.4: Visualization of the dot product between hidden states and $\mathbf{d}_{\text{safe}}$ across layers and token positions on Llama-3-70B-it.
Figure A.5: Visualization of the dot product between hidden states and $\mathbf{d}_{\text{safe}}$ across layers and token positions on Phi-3-mini-it.
Figure A.6: Visualization of the dot product between hidden states and $\mathbf{d}_{\text{safe}}$ across layers and token positions on Qwen2-7B-it.
@@ A.2.5 Additional Results on Scaling Factor $\alpha$
Table A.4: Performance against adaptive jailbreak attacks. Results shown as Baseline / Antidote DSR (%).
Model GCG PAIR AutoDAN
Llama-2-7B-chat 46 / 83 78 / 91 67 / 86
Llama-2-13B-chat 70 / 86 85 / 93 76 / 87
Llama-3-8B-it 73 / 93 89 / 95 78 / 91
A.2.6 Comparison of DSR and Length Controlled Win Rate
We provide a detailed comparison of Length Controlled Win Rate (Win Ratelc) across different models and defense methods. As shown in Table A.5, the differences in Win Ratelc across various methods remain relatively small compared to the non-length-controlled results in Table 1. However, our proposed method, Jailbreak Antidote, consistently achieves higher Win Ratelc in this controlled setting. This improvement likely stems from the conservative nature of our defense strategy, which generates fewer but more aligned responses, thereby maintaining both safety and effectiveness under length-controlled conditions.
Table A.5: Comparison of Defense Success Rate (DSR) and Length Controlled Win Rate (Win Ratelc) across different models and defense methods. The best, second and third scores are highlighted.
Model Safety -Utility
Baseline
In-Context Learning
Paraphrase
Perplexity Filter
Self Reminder
Semantic Smooth LLM
Smooth LLM
Jailbreak Antidote
Gemma-2-2B-it DSR $\uparrow$ 29.2 54.1 69.7 30.5 36.1 74.5 46.5 71.8
Win Ratelc $\uparrow$ 50.0 44.9 35.8 50.6 47.7 31.6 31.4 52.3
Phi-3-mini-it DSR $\uparrow$ 53.2 55.4 75.5 54.9 55.4 70.5 71.3 79.6
Win Ratelc $\uparrow$ 50.0 42.6 36.4 47.6 44.1 27.8 20.1 52.2
Qwen-1.5-7B-it DSR $\uparrow$ 29.2 54.1 69.7 30.5 36.1 74.5 46.5 71.8
Win Ratelc $\uparrow$ 50.0 44.6 35.5 50.4 47.6 31.7 31.6 52.3
Qwen-2-7B-it DSR $\uparrow$ 55.3 57.6 70.1 57.3 67.4 81.9 60.4 95.5
Win Ratelc $\uparrow$ 50.0 37.5 35.4 51.7 49.6 33.1 32.7 52.1
Llama-3-8B-it DSR $\uparrow$ 68.9 71.7 79.0 67.9 78.9 88.1 84.2 99.4
Win Ratelc $\uparrow$ 50.0 38.6 35.4 51.6 39.6 31.9 32.4 52.8
Llama-3.1-8B-it DSR $\uparrow$ 63.1 56.2 72.1 64.0 68.6 86.6 69.0 78.0
Win Ratelc $\uparrow$ 50.0 36.5 32.4 51.3 41.6 26.1 32.7 52.3
Gemma-2-9B-it DSR $\uparrow$ 54.5 56.7 75.8 55.1 63.1 79.4 46.5 78.1
Win Ratelc $\uparrow$ 50.0 38.7 31.6 51.1 42.1 33.6 32.6 47.8
Llama-3-70B-it DSR $\uparrow$ 61.4 61.8 76.1 61.6 71.8 83.9 88.2 100
Win Ratelc $\uparrow$ 50.0 36.4 35.6 50.3 42.5 33.8 35.6 53.6
Qwen-2-72B-it DSR $\uparrow$ 62.7 61.5 65.0 65.2 71.0 72.3 69.8 93.9
Win Ratelc $\uparrow$ 50.0 35.6 34.4 48.7 45.8 30.9 33.7 53.2
A.2.7 Evaluation Against Adversarial Attacks
We evaluated Jailbreak Antidote against three representative adversarial attack strategies: Gradient-based Content Generation (GCG) (Zou et al., 2023b), PAIR (Chao et al., 2023), and AutoDAN (Liu et al., 2024). In these experiments, we fixed the parameters of Jailbreak Antidote (e.g., $\mathbf{d}_{safe}^{l}$, $\mathbf{m}^{l}$, and $\alpha$) and allowed the attack methods to dynamically generate prompts aimed at bypassing the defense. This approach directly tests the robustness and efficiency of our method when faced with adversarial strategies targeting the model’s defenses. The results are summarized in Table A.4.
Our findings indicate that Jailbreak Antidote significantly improves resilience to such attacks across all tested models and strategies. For example, on Llama-2-7B-chat, the defense success rate against GCG increased from 46% to 83% when Jailbreak Antidote was applied. These results validate the robustness and versatility of our approach, even when adversarial attacks dynamically attempt to circumvent the defense within the constraints of the fixed parameters.
A.2.8 Analysis of False Positive Rate in Safety Blocking
Evaluating the false positive rate of safety blocking in language models is inherently challenging due to the open-ended nature of generative tasks. While the DSR captures the model’s ability to block harmful prompts, and the Win Rate measures its utility on benign tasks, quantifying how often the model incorrectly refuses benign queries (false positives) requires additional analysis.
To estimate the false positive rate, we conducted a simple evaluation based on the responses generated in AlpacaEval. Specifically, we defined a response as a clear refusal if it begins with phrases like ”I cannot” or similar expressions indicating refusal (Zou et al., 2023b). Using this heuristic, we calculated the refusal rate for varying values of the scaling factor $\alpha$.
Figure A.9 shows the relationship between the refusal rate and the Win Rate for the Llama-3-8B-it and Llama-3-70B-it models. The results indicate a clear trade-off: as $\alpha$ increases, the refusal rate rises, and the Win Rate correspondingly decreases. This confirms that larger $\alpha$ values make the model more conservative, leading to a higher likelihood of rejecting borderline benign prompts.
Figure A.9: Relationship between refusal rate and Win Rate for (a) Llama-3-8B-it and (b) Llama-3-70B-it across varying $\alpha$ values. The refusal rate increases as $\alpha$ grows, resulting in a decline in Win Rate.
Figure A.10: Defense Tokens versus DSR for different defense methods across various models. Each point represents a defense method, with the x-axis showing the number of defense tokens and the y-axis showing the DSR.
Model Pre-Inference Time
Gemma-2-2B-it 39.6s
Phi-3-mini-it 43.2s
Qwen-2-7B-it 48.2s
Llama-3-8B-it 54.3s
Gemma-2-9B-it 1m26.4s
Llama-3-70B-it 5m13.3s
Qwen-2-72B-it 5m32.5s
Table A.6: Pre-inference time for various models.
Model JailbreakBench AlpacaEval
Gemma-2-2B-it 0.96x 0.98x
Phi-3-mini-it 0.94x 0.97x
Qwen-2-7B-it 0.85x 0.98x
Llama-3-8B-it 0.82x 1.02x
Gemma-2-9B-it 0.98x 1.04x
Llama-3-70B-it 0.79x 1.02x
Qwen-2-72B-it 0.86x 1.01x
Table A.7: Inference time relative to the baseline. Values are averaged across all attack methods online JailbreakBench.
A.2.9 Comparative Analysis of AlpacaEval with MMLU and HellaSwag
To further substantiate the utility of AlpacaEval, we compare its results with those of two downstream tasks, MMLU (Hendrycks et al., 2021) and HellaSwag (Zellers et al., 2019), on the Llama-3-8B-it model, as shown in Table A.8. These results provide a more comprehensive view of model performance across diverse tasks, including factual knowledge (MMLU) and commonsense reasoning (HellaSwag).
Table A.8: Comparison of AlpacaEval (Win Rate), MMLU, and HellaSwag on Llama-3-8B-it.
Safety -Utility
Baseline
In-Context Learning
Paraphrase
Perplexity Filter
Self Reminder
Semantic Smooth LLM
Smooth LLM
Jailbreak Antidote
Win Rate 50.0 38.9 35.5 52.2 39.4 31.8 32.4 53.0
MMLU 66.7 65.3 64.3 66.7 65.8 62.8 63.2 67.4
HellaSwag 82.2 80.1 79.5 82.3 78.6 80.6 73.9 82.5
The results highlight a consistent trend where methods demonstrating higher utility in AlpacaEval also perform competitively on downstream benchmarks. Notably, Jailbreak Antidote achieves the best performance across all three metrics, demonstrating its robustness. Meanwhile, other methods show varying degrees of alignment across the tasks, which underscores the importance of multi-metric evaluations to capture diverse aspects of model behavior.